I need to understand how to solve this equation:
College Students spend $183 more each year on textbooks than on computer equipment. They spend a total of $819 on textbooks and computer equipment each year. How much is spent each year on textbooks and how much is spend on computer equipment.

Answer by checkley75(3666)

LET X=AMOUNT SPENT ON COMPUTERS
THEN (X+183)=AMOUNT SPENT ON BOOKS.
TOTAL SPENT=819 THEN
X+(X+183)=819
2X+183=819
2X=819-183
2X=636
X=636/2
X=318 ANSWER FOR THE AMOUNT SPENT ON COMPUTERS.
PROOF
318+(318+183)=819
318+501=819
819=819